# A Quantitative Evaluation of the Expressivity of BMI, Pose and Gender in Body Embeddings for Recognition and Identification

### v1

@@ I INTRODUCTION
Deep learning models are primarily designed to learn specific target attributes during training; however, they frequently encode unintended, image-related attributes that can introduce significant biases into the system. In the domain of biometrics, particularly face recognition, Hill et al. [18] demonstrated that deep networks form identity representations that inherently cluster based on gender. Moreover, these identity embeddings have been shown to encode other latent characteristics such as pose, age, and lighting conditions [18, 29, 34]. The presence of these latent attributes can significantly influence the accuracy of recognition algorithms, as they may inadvertently affect model predictions [15, 26]. Evaluating these biases or understanding the correlation between attributes and network features require a systematic analysis of how these attributes are embedded and how they influence model behavior. A deeper understanding of these phenomena necessitates investigating how facial or body attributes are encoded in identity representations and how they shape predictive outcomes. In this context, Dhar et al. [12] introduced the concept of expressivity, for face recognition, a metric that quantifies the relationship between learned network features and specific attributes, thereby enhancing the interpretability of face recognition models. Building on this concept, we extend the framework of expressivity to the domain of person re-identification (ReID) with the goal of evaluating how body-related features are embedded within ReID models trained primarily for identity recognition.
@@
Person re-identification (ReID) is a well-established research area with a range of real-world applications, including smart city infrastructure for public safety and traffic management [2, 22] and autonomous driving systems for pedestrian detection and tracking [4, 41]. The primary objective of ReID is to accurately match and retrieve pedestrian identities across non-overlapping camera views, varying time frames, and distinct locations, all while addressing challenges such as pose variations, appearance diversity, and environmental conditions [16, 17, 46]. Significant progress has been achieved in improving re-identification accuracy through the development of deep learning methods, which can broadly be categorized into image-based and video-based approaches. Image-based ReID methods focus on selecting the most distinctive frame and extracting fine-grained spatial features, while video-based approaches aggregate temporal information across multiple frames to produce more robust identity representations. Recent advancements have increasingly combined these approaches, leveraging the strengths of both image-level detail and temporal consistency to achieve state-of-the-art performance. Despite these advancements, most deep learning-based ReID systems are trained to identify individuals based on visual body features, without explicitly learning specific body-related attributes. These models generate identity representations derived from body cues; however, similar to face recognition systems, ReID networks often unintentionally encode additional attributes related to body characteristics. To address this, our work conducts a comprehensive analysis of the attributes correlated with feature embeddings generated by the state-of-the-art SemReID model. As ReID systems are increasingly deployed in real-world applications, there is a growing demand for explainable and transparent models. Understanding how various attributes are encoded across internal network layers is crucial for interpreting identity predictions and identifying potential sources of algorithmic bias. The following are the conceptual and experimental contributions of our paper:
@@
  • We present the first investigation into the encoding of body attributes within the layers of a large-scale Vision Transformer (ViT)-based foundation model for person ReID. To enhance the interpretability of large-scale deployable ReID systems, we propose a novel post-hoc framework that explains how internal representations influence identity predictions. This achievement underscores the robustness of our approach, despite the inherent complexity of the model and the diversity of the dataset.
@@
  • In the final attention layer of the SemReID network, we observe the following order of expressivity for body attributes: BMI $>$ Pitch $>$ Yaw $>$ Gender. This ranking highlights the varying degrees of influence that different attributes have on the network’s predictions.
@@
II RELATED WORKS
Person re-identification (ReID) is a critical task in computer vision that focuses on matching individuals across non-overlapping camera views, often under varying conditions such as illumination, clothing, pose, and occlusion [16, 17, 21, 46]. Numerous works have been proposed to address this challenge in various ReID domains, such as Clothes-Changing ReID (CC-ReID)[16], video ReID[5, 19, 43, 45, 42], unconstrained ReID[11, 27, 31, 30, 47], and short-term ReID[9, 45, 40, 48]. Among these methods, the SemReID [20] approach achieves state-of-the-art (SoTA) performance across all four domains. Though the interpretability of these specific models have not been extensively explored in prior literature, several works have examined the explainability of recognition algorithms in general, primarily focusing on face recognition.
The study of bias and interpretability in biometrics has long been a subject of significant interest [36, 13, 38, 32, 33]. Schumann et al. [35] introduced an auxiliary network to enhance the performance of ReID systems by incorporating complementary information into CNN features. More recently, Myers et al. [28] demonstrated that linguistic and non-linguistic representations of body shape can provide complementary identity information, improving identification in specific applications. These methods rely on the change in predictions with respect to a concept or attribute to interpret a network’s sensitivity to the attribute. Yin et al. [44] introduced a spatial activation diversity loss to maintain interpretability in face recognition networks during training. Similarly, Kim et al. [23] proposed a generative approach utilizing representative exemplars (prototypes) to enhance model interpretability. However, as noted in [24], these methods are limited to models trained from scratch and cannot be applied to pre-trained networks or those already deployed in practice.
Post-hoc interpretability methods have also been explored to understand trained models, with TCAV (Testing with Concept Activation Vectors) [24] being one of the most influential approaches in this domain. TCAV interprets a model’s behavior by analyzing its sensitivity to user-defined concepts and achieves this by learning Concept Activation Vectors (CAVs) through training a linear classifier to distinguish activations produced by concept examples. While TCAV is effective for discrete physical concepts like color or texture, it is not suitable for assessing sensitivity to continuous or abstract attributes, such as pose angle or BMI. This limitation arises because generating negative examples (images without the concept) is challenging for omnipresent and continuous attributes. Furthermore, TCAV requires test images to belong to one of the training classes, as its computations rely on changes in logits for specific classes, rendering it incompatible with scenarios involving unseen subjects or faces. Other post-hoc methods, such as [1], utilize linear classifiers to analyze intermediate network layers, while Koh and Liang [25] proposed influence functions to measure model sensitivity to infinitesimal perturbations in training data. However, these methods are unsuitable for evaluating sensitivity to physical attributes like pose or orientation, as they primarily focus on localized perturbations. Saliency- and attention-based approaches [37, 8] offer another class of interpretability methods, generating attention maps to visualize spatial regions influencing model predictions. For person ReID, Chen et al. [10] introduced a learnable, pluggable interpreter for CNN-based models, which decomposes image-pair distances into attribute-based contributions and visualizes attention maps for discriminative attributes. While effective, this method relies on attribute-guided metric distillation, involves computationally intensive post-hoc operations, and is not adapted for transformer-based backbones. Moreover, while saliency maps can highlight spatial regions, they cannot assess sensitivity to abstract or non-localized concepts, such as BMI or pitch/yaw. Few studies explicitly address the interpretability of face recognition models on the basis of attribute hierarchy, for example Hill et al. [18] demonstrated a hierarchy within facial feature representations, showing that identity is nested under sex, illumination is nested under identity, and viewpoint is nested under illumination. Dhar et al. [12] further extended this analysis by conducting an expressivity-based evaluation of facial attributes, identifying a hierarchy where age has the greatest influence, followed by sex and yaw.
@@ II RELATED WORKS
We propose the use of expressivity as a measure for person ReID systems, quantifying the predictability of an attribute within a given set of features extracted using the model. Unlike prior methods, expressivity can be computed for both categorical and continuous attributes, allowing for a direct comparison of attribute predictability. Additionally, expressivity can be calculated agnostic to the ReID network backbone. In this work, we demonstrate results using a SoTA Vision Transformer (ViT)-based ReID model. This framework provides valuable insights into how body-related features are embedded within ReID models and their influence on performance, laying the groundwork for more interpretable and explainable ReID systems.
III Proposed Method
@@
Our approach attempts to find the correlations between the learnt features by a state-of-the-art (SoTA) body recognition model and sensitive attributes, as highlighted in Figures 1 and 2. The predictability of attributes from a given set of body descriptors reflects the amount of attribute-relevant information encoded within those descriptors. To quantify this information, we employ Mutual Information (MI) as shown in Equation 1. MI is a fundamental quantity for measuring the relationship between random variables, indicating how much knowledge of one variable reduces uncertainty about the other. By estimating the MI between features learned by the body recognition model and their corresponding sensitive attributes, we assess the degree to which these descriptors encode attribute information. Since MI captures non-linear statistical dependencies between variables and is applicable to both categorical and continuous attributes, this approach provides a unified and consistent measure across attribute types. To develop a general-purpose estimator, we utilize the widely recognized formulation of MI as the Kullback-Leibler (KL) divergence (Kullback, 1997) between the joint distribution and the product of the marginal distributions of two random variables $X$ and $Z$, as expressed in Equation 2.
@@
III-A Problem Setup
Figure 1: Integrating the MINE block in the ViT based SemReID [20] backbone to compute the expressivity of features with respect to attributes such as BMI, gender, pitch and yaw. The internal structure of the MINE block is shown in the next figure.
Figure 2: The internal structure of the MINE block employs a simple MLP with two hidden layers to compute the expressivity of $m$-dimensional features $F$. By augmenting these features with an attribute vector $A$, the input to the network is extended to $(m+1)$-dimensions.
Our dataset comprises body images of different individuals captured under varying conditions and at different distances. Each image is annotated with an identity label and several sensitive attributes, including gender ($g$), height ($h$), weight ($w$), body mass index ($BMI$), which is computed from $h$ and $w$, as well as pitch angles ($p$) and yaw angles ($y$). These attributes collectively form a diverse set of information, enabling a comprehensive analysis of how sensitive attributes are encoded in the learned features. We denote the set of learned feature descriptors as $\mathbf{F}$, and the corresponding sensitive attributes as $A$. The primary objective of this analysis is to quantify and explore the correlations between $\mathbf{F}$ and $A$ using Mutual Information Neural Estimation (MINE). Specifically, we aim to estimate the MI, denoted as $I_{\theta}(\mathbf{F},\mathbf{A})$, to gain insights into how effectively the learned features capture attribute-relevant information.
To achieve this, each image $x_{i}$ undergoes a series of preprocessing steps before being passed through the SemReID model. The model extracts feature descriptors $f_{i}$, where $f_{i}\in\mathbb{R}^{m}$, representing the encoded identity and attribute information for each image. These descriptors are then concatenated to form the feature matrix $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$, where $F\in\mathbb{R}^{n\times m}$. The sensitive attribute vector $A\in\mathbb{R}^{n\times 1}$, containing information such as gender, pose, and identity, is then combined with $F$ to form an augmented matrix $\mathbf{X}=[\mathbf{F}|\mathbf{A}]$. This augmented matrix $X$ is subsequently used by the MINE network to estimate the MI between $\mathbf{F}$ and $\mathbf{A}$.
@@ III-A Problem Setup
III-B Attributes and Their Relevance
@@
We compute the expressivity of four annotated attributes: $g$, $BMI$, $p$ and $y$ in the extracted features. In Figure 3, we verify that the dataset we utilize shows enough variation with respect to these attributes, so that we can ensure that expressivity (which is a lower bound estimate of MI) is an accurate model for the corresponding attributes. When considering g, the vector $\mathbf{A}$ is a discrete vector having a value of 1 if the gender is male and 0 if female while for $BMI$, y and p (in degrees) the values of the vector values are continuous. These attributes play a vital role in person ReID tasks, as they influence the expressivity of the model’s learned features. Attributes like $BMI$ or pose are particularly challenging to disentangle from identity for this task, making them ideal for analyzing feature relevance.
III-C Expressivity of Body Features
@@
Understanding the expressivity of learned features in deep networks is critical for tasks that rely on nuanced feature representations, such as person re-identification (ReID). Tishby and Zaslavsky [39] introduced the concept of utilizing MI as a quantitative metric to assess how well information is retained or transformed across the layers of a deep network. MI measures the dependency between random variables, offering insights into the trade-offs between compression and informativeness at various stages of a network. By quantifying MI, one can directly evaluate how effectively the network balances these competing objectives.
@@
However, estimating MI for high-dimensional continuous variables is computationally challenging due to the need to compute probability density functions of the underlying distributions. Traditional methods often rely on discretization or kernel density estimation, both of which suffer from scalability issues as dimensionality increases. To overcome this, Belghazi et al. [3] proposed Mutual Information Neural Estimation (MINE), a scalable framework that approximates MI using a neural network. This approach bypasses the need for explicit density computations by optimizing a neural network-based lower bound of MI, making it suitable for high-dimensional and complex datasets.
The MI between learned feature descriptors $\mathbf{F}$ and sensitive attributes $\mathbf{A}$ is a crucial metric in evaluating the expressivity of the learned features. In the context of this work, $\mathbf{F}$ represents the feature embeddings produced by the SemReID model, while $\mathbf{A}$ denotes associated sensitive attributes such as gender, pose, and identity. The MI is mathematically defined as:
$I_{\theta}(\mathbf{F},\mathbf{A})=\sup_{\theta\in\Theta}\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]-\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right],$ (3)
@@
where $T_{\theta}(f,a)$ is a neural network parameterized by $\theta$, designed to approximate the MI. The joint expectation $\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]$ measures the network’s output when conditioned on the true joint distribution of features and attributes. In contrast, the term $\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$ normalizes the MI estimate to ensure it captures only the dependency between $\mathbf{F}$ and $\mathbf{A}$, excluding any bias from their marginal distributions.
Computational Steps for Mutual Information Estimation
@@
The MI lower bound is approximated as the difference between the joint and marginal expectations, yielding the following objective function:
@@
In the context of person ReID, this framework is particularly valuable for understanding the expressivity of the feature descriptors generated by the model. The neural network $T_{\theta}$ is trained to approximate the MI between the learned features $\mathbf{F}$ and sensitive attributes $\mathbf{A}$. By iteratively computing joint and marginal expectations and updating $\theta$, the MI provides a robust metric to quantify how much attribute-relevant information is encoded in the features. At convergence, the MI value reflects the extent to which the model’s representations capture sensitive attribute information, offering insights into the expressivity and fairness of the learned features.

### v3

@@ 1 Introduction
Deep learning models are trained to learn specific target attributes, but often encode unintended image-related attributes that can adversely affect model performance and fairness. In the domain of biometrics, particularly face recognition, Hill et al. [18] demonstrated that deep networks form identity representations that inherently cluster based on gender. Moreover, these identity embeddings have been shown to encode other latent characteristics such as pose, age, and lighting conditions [18, 31, 36]. The presence of these latent attributes can significantly influence the accuracy of recognition algorithms, as they may inadvertently affect model predictions [15, 26]. Evaluating these biases or understanding the correlation between attributes and network features require a systematic analysis of how these attributes are embedded and how they influence model behavior. A deeper understanding of these phenomena necessitates investigating how facial or body attributes are encoded in identity representations and how they shape predictive outcomes. In this context, Dhar et al. [12] introduced the concept of expressivity, for face recognition, a metric that quantifies the relationship between learned network features and specific attributes, thereby enhancing the interpretability of face recognition models. Building on this concept, we extend the framework of expressivity to the domain of person ReID with the goal of evaluating how body-related features are embedded within ReID models trained primarily for recognition.
@@
Person ReID is a well-established research area with a range of real-world applications, including smart city infrastructure for public safety and traffic management [2, 22] and autonomous driving systems for pedestrian detection and tracking [4, 44]. The primary objective of ReID is to accurately match and retrieve pedestrian identities across non-overlapping camera views, varying time frames, and distinct locations, while addressing challenges such as pose variations, appearance diversity, and environmental conditions [16, 17, 49]. Significant progress has been achieved in improving ReID accuracy through the development of deep learning methods, which can broadly be categorized into image-based and video-based approaches. Image-based ReID methods focus on selecting the most distinctive frame and extracting fine-grained spatial features, while video-based approaches aggregate temporal information across multiple frames to produce more robust identity representations. Recent advancements have increasingly combined these approaches, leveraging the strengths of both image-level detail and temporal consistency to achieve state-of-the-art (SoTA) performance. Despite these advancements, most deep learning-based ReID systems are trained to identify individuals based on visual body features, without explicitly learning specific body-related attributes. These models generate identity representations derived from body cues; however, similar to face recognition systems, ReID networks often unintentionally encode additional attributes related to body characteristics. To address this, our work conducts a comprehensive analysis of the attributes correlated with feature embeddings generated by the three SoTA ReID models and compares their robustness. Recently, we have come across the work of Metz et al.[29], where they also attempt to identify what information beyond identity is stored in the feature vectors from learned body recognition models. While they used an empirical method by training a logistic regression model to predict gender from image embeddings, it primarily demonstrates the presence of linearly separable attribute information. This method relies on performance metrics from a downstream classifier and does not capture the underlying statistical dependencies among attributes and representations. We adopt an information-theoretic perspective by applying Mutual Information Neural Estimation (MINE) to directly quantify the dependency among attribute variables and deep body recognition features. This allows us to measure how much information about an attribute is encoded in the feature space, regardless of classifier performance. Thus by moving beyond specific prediction and directly analyzing feature–attribute dependencies, our method offers a more reliable and theoretically grounded evaluation of attribute leakage and representational bias understanding. As ReID systems are increasingly deployed in real-world applications, there is a growing demand for explainable and transparent models. Understanding how various attributes are encoded across internal network layers is crucial for interpreting identity predictions and identifying potential sources of algorithmic bias. The following are the conceptual and experimental contributions of our paper:
@@
  • We present the first investigation into the encoding of body attributes within the layers of a large-scale Vision Transformer (ViT)-based self-supervised foundation model called semReID[20]. We also compare our results against two other transformer based benchmark models. To enhance the interpretability of large-scale deployable ReID systems, we present a post-hoc framework that explains how internal representations influence identity predictions. This underscores the robustness of our method, despite the inherent complexity of the model and the diversity of the dataset.
@@
  • In the final attention layer of the SemReID network as well as the other networks, we observe the following order of expressivity for body attributes: BMI $>$ Pitch $>$ Gender$>$ Yaw. This ranking highlights the varying degrees of influence that different attributes have on the network’s predictions.
@@
2 Related Works
Person ReID aims to match individuals across non-overlapping camera views under challenging conditions such as illumination, clothing, pose, and occlusion [16, 17, 21, 49]. Extensive efforts have addressed this problem across domains like Clothes-Changing ReID (CC-ReID) [16], video ReID [5, 19, 46, 48, 45], unconstrained ReID [11, 28, 33, 32, 50], and short-term ReID [9, 48, 42, 51]. Among these, SemReID [20] achieves SoTA performance across all four domains. While ReID interpretability remains underexplored, broader recognition systems, especially face recognition have received more attention.
Bias and interpretability in biometrics have long been studied [38, 13, 40, 34, 35]. Schumann et al. [37] used an auxiliary network to enrich CNN features, and Myers et al. [30] leveraged both linguistic and non-linguistic body representations for identity prediction. These works analyze model sensitivity to attributes via concept-based prediction changes. Yin et al. [47] introduced a spatial activation diversity loss to preserve interpretability in face recognition, while Kim et al. [23] proposed a prototype-based generative model. However, as noted in [24], such methods are limited to models trained from scratch and do not generalize to deployed networks. Post-hoc interpretability methods offer alternatives, notably TCAV [24], which measures sensitivity to user-defined concepts via Concept Activation Vectors (CAVs) learned through linear classification. While effective for discrete attributes like color or texture, TCAV struggles with continuous or omnipresent attributes (e.g., BMI, pose), where defining negative examples is difficult. TCAV also requires test images to belong to seen classes, limiting use in open-set scenarios. Other methods include layer-wise linear probes [1], influence functions [25], and saliency-based approaches [39, 8]. For ReID specifically, Chen et al. [10] proposed a pluggable interpreter that attributes image-pair distances to visual cues but depends on metric distillation and is tailored to CNNs. Saliency maps, while helpful for spatial focus, cannot explain abstract or non-localized attributes. Studies in face recognition have further examined attribute hierarchies. Hill et al. [18] revealed that identity representations are nested under sex, illumination, and viewpoint, while Dhar et al. [12] used expressivity-based evaluations to identify a hierarchy where age dominates, followed by sex and yaw.
@@ 2 Related Works
We propose expressivity as a general framework to assess person ReID systems by quantifying how well an attribute can be predicted from learned features. Unlike prior approaches, expressivity applies to both categorical and continuous attributes and is agnostic to model backbone. We demonstrate its utility using ViT-based ReID models, offering insights into how body-related features are embedded and their impact on model performance, paving the way for more interpretable and explainable ReID systems.
Figure 1: Integrating the MINE block with the ViT based SemReID [20] backbone to compute the expressivity of features with respect to attributes such as BMI, gender, pitch and yaw. The internal structure of the MINE block employs a simple MLP with two hidden layers to compute the expressivity of $m$-dimensional features $F$. By augmenting these features with an attribute vector $A$, the input to the network is extended to $(m+1)$-dimensions. All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures.
3 Proposed Method
@@
Our method as seen in Figure 1 tries to find the correlations between the learnt features by body recognition models and attributes. The predictability of attributes from a given set of body descriptors reflects the amount of attribute-relevant information encoded within those descriptors. To quantify this information, we employ Mutual Information (MI) as shown in Equation 1. MI is a fundamental quantity for measuring the relationship between random variables, indicating how much knowledge of one variable reduces uncertainty about the other. By estimating the MI between features learned by the body recognition model and their corresponding sensitive attributes, we assess the degree to which these descriptors encode attribute information. Since MI captures non-linear statistical dependencies between variables and is applicable to both categorical and continuous attributes, this approach provides a unified and consistent measure across attribute types. To develop a general-purpose estimator, we utilize the widely recognized formulation of MI as the Kullback-Leibler (KL) divergence (Kullback, 1997) between the joint distribution and the product of the marginal distributions of two random variables $X$ and $Z$, as expressed in Equation 2.
@@
3.1 Problem Setup
Our dataset comprises body images of different individuals captured under varying conditions and at different distances. Each image is annotated with an identity label and several sensitive attributes, including gender ($g$), height ($h$), weight ($w$), body mass index ($BMI$), which is computed from $h$ and $w$, as well as pitch angles ($p$) and yaw angles ($y$). These attributes collectively form a diverse set of information, enabling a comprehensive analysis of how sensitive attributes are encoded in the learned features. We denote the set of learned feature descriptors as $\mathbf{F}$, and the corresponding sensitive attributes as $A$. The primary objective of this analysis is to quantify and explore the correlations between $\mathbf{F}$ and $A$ using MINE. Specifically, we aim to estimate the MI, denoted as $I_{\theta}(\mathbf{F},\mathbf{A})$, to gain insights into how effectively the learned features capture attribute-relevant information. To achieve this, each image $x_{i}$ undergoes a series of preprocessing steps before being passed through the person ReID models. The model extracts feature descriptors $f_{i}$, where $f_{i}\in\mathbb{R}^{m}$, representing the encoded identity and attribute information for each image. These descriptors are then concatenated to form the feature matrix $\mathbf{F}=[f_{1},f_{2},\dots,f_{n}]^{T}$, where $F\in\mathbb{R}^{n\times m}$. The sensitive attribute vector $A\in\mathbb{R}^{n\times 1}$, containing information such as gender, pose, and identity, is then combined with $F$ to form an augmented matrix $\mathbf{X}=[\mathbf{F}|\mathbf{A}]$. This augmented matrix $X$ is subsequently used by the MINE network to estimate the MI between $\mathbf{F}$ and $\mathbf{A}$.
@@ 3.1 Problem Setup
3.2 Attributes and Their Relevance
@@
We compute the expressivity of four annotated attributes: $g$, $BMI$, $p$ and $y$ in the extracted features. In Figure  2, we verify that the dataset we utilize shows enough variation with respect to these attributes, so that we can ensure that expressivity (which is a lower bound estimate of MI) is an accurate model for the corresponding attributes. When considering g, the vector $\mathbf{A}$ is a discrete vector having a value of 1 if the gender is male and 0 if female while for $BMI$, y and p (in degrees) the values of the vector values are continuous. These attributes play a vital role in person ReID tasks, as they influence the model’s learned features.
3.3 Expressivity of Body Features
@@
Understanding the expressivity of learned features in deep networks is critical for tasks that rely on nuanced feature representations, such as person ReID. Tishby and Zaslavsky [41] introduced the concept of utilizing MI as a quantitative metric to assess how well information is retained or transformed across the layers of a deep network. MI measures the dependency between random variables, offering insights into the trade-offs between compression and informativeness at various stages of a network. By quantifying MI, one can directly evaluate how effectively the network balances these competing objectives.
@@
However, estimating MI for high-dimensional continuous variables is computationally challenging due to the need to compute probability density functions of the underlying distributions. Traditional methods often rely on discretization or kernel density estimation, both of which suffer from scalability issues as dimensionality increases. To overcome this, Belghazi et al. [3] proposed MINE, a scalable framework that approximates MI using a neural network. This bypasses the need for explicit density computations by optimizing a neural network-based lower bound of MI, making it suitable for high-dimensional and complex datasets.
@@
The MI between learned feature descriptors $\mathbf{F}$ and sensitive attributes $\mathbf{A}$ is a crucial metric in evaluating the expressivity of the learned features. In the context of this work, $\mathbf{F}$ represents the feature embeddings produced by the ReID model, while $\mathbf{A}$ denotes associated sensitive attributes such as gender, pose, and identity. The MI approximate is mathematically defined as $I_{\theta}(\mathbf{F},\mathbf{A})=\sup_{\theta\in\Theta}\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]-\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$, where $T_{\theta}(f,a)$ is a neural network parameterized by $\theta$, designed to approximate the MI. The joint expectation $\mathbb{E}_{P_{FA}}\left[T_{\theta}(f,a)\right]$ measures the network’s output when conditioned on the true joint distribution of features and attributes. In contrast, the term $\log\mathbb{E}_{P_{F}\otimes P_{A}}\left[e^{T_{\theta}(f)}\right]$ normalizes the MI estimate to ensure it captures only the dependency between $\mathbf{F}$ and $\mathbf{A}$, excluding any bias from their marginal distributions.
Computational Steps for MI Estimation:
@@
The MI lower bound is approximated as the difference between the joint and marginal expectations, yielding the objective function:
@@
In the context of person ReID, this framework is particularly valuable for understanding the expressivity of the feature descriptors generated by the model. The neural network $T_{\theta}$ is trained to approximate the MI between the learned features $\mathbf{F}$ and sensitive attributes $\mathbf{A}$. By iteratively computing joint and marginal expectations and updating $\theta$, the MI provides a robust metric to quantify how much attribute-relevant information is encoded in the features. At convergence, it reflects the extent to which the model’s representations capture sensitive attribute information, offering insights into the expressivity and fairness of the learned features.
4 Experiments
4.1 Dataset and Settings
Figure 3: Attribute annotated exemplar images from the BRIAR dataset. All subjects involved provided informed consent for their participation, including the use of their images in research publications and figures.
We use the BRIAR 1–5 dataset [11], a large-scale unconstrained person re-identification benchmark comprising over 1 million images and 40,000 videos captured under real-world conditions, including varying clothing, distances (100m–1km), altitudes (e.g., UAV), and environmental challenges like occlusion, blur, and turbulence. BRIAR includes five progressively complex subsets (BRIAR-1 to 5), increasing in identities, distractors, and capture variability. For our study, we extract 704,999 frames from 382,229 images and 170,522 videos, covering 2,077 unique identities (887 male and 1,190 female subjects). Figure 3 shows examples of images and attribute annotations from our curated subset.
4.2 Integration of MINE with ReID Models